This year Evan graduated from college and took a job as a deliveryman in the city. Evan was paid a salary of $74,700 and he received $700 in hourly pay for part-time work over the weekends. Evan summarized his expenses as follows:

Cost of moving his possessions to the city (125 miles away) $ 1,200
Interest paid on accumulated student loans 2,880
Cost of purchasing a delivery uniform 1,480
Contribution to State University deliveryman program 1,340

Calculate Evan's AGI and taxable income if he files single. Assume that interest payments were initially required on Evan’s student loans this year.

Evan's AGI : ?

Taxable Income: ?
Answer #1

Computation of Evans AGI and Taxable Income

Particulars $
Salary income 74,700
Part time income 700
Modified AGI for purpose of student loan 75,400
Student Loan deduction (1600)
AGI 73,800
Less standard deduction (12,200)
Taxable income 61,600

Evans modified AGI for purpose of student loan exceeds $70,000 which is threshold limit. Exceeds by an amount of (75400-70000) = 5400. So applicable ratio is 5400/15000*100 = 36% . His deduction education loan is subject to phase out .

Deduction allowed = $2500 - 2500*36% = 2500-900 = $1600

Standard deduction allowable is $12,200 for individual tax payers.